OpenStudy (anonymous):

can someone please help make since of absolute value? The problem is |x-8|-2=-2. I figured out the answer to one part but do not understand if I answered it correctly. Can someone please solve and tell me how you got the answer? Thanks in advance

OpenStudy (anonymous):

|x-8|-2=-2. ==>|x-8| = 0 ===> x-8=0 ==> x=8

OpenStudy (anonymous):

|x-8|=0 means the x values that are exactly 0 units from 8. the only number that is exactly 0 units from 8 is 8.